OpenStudy (anonymous):

Help me find the slope and Y intercept? Y = -3/2x + 4 –3x + 2y = 8

OpenStudy (midhun.madhu1987):

The slope-intercept form of a line is given by: y = mx + c where m=slope and c = y-intercept The given line is of the form Y = -3/2x + 4 So what is the value of m and c?

OpenStudy (anonymous):

3/2 is m and 4 is Y?

OpenStudy (midhun.madhu1987):

check for the sign of m tooo

OpenStudy (anonymous):

@calculusfunctions Help? I need the help.

OpenStudy (calculusfunctions):

Yes @RedBlade ? How can I help you? @midhun.madhu1987 explained the slope-intercept method perfectly. To that I will add that if given the equation of a line in the standard (Cartesian) form Ax + By + C = 0, then the slope m = -A/B, the y-intercept is -C/B, and the x-intercept is -C/A